Solve the Bessel's equation of order zero

y''(x)+(1/x)y'(x)+y(x)=0

With boundary conditions y'(0)=0, y(1)=1


1
Expert's answer
2021-04-15T07:39:40-0400

Multiplying on x2

x2y"(x)+xy'(x)+x2y(x)=0

"y(x)=\sum_{n=0}^{\propto}a_nx^{r+n}" , for a not equal 0, x>0

"y'(x)=\sum_{n=0}^{\propto}a_n(r+n)x^{r+n-1}"

"y"(x)=\sum_{n=0}^{\propto}a_n(r+n)(r+n-1)x^{r+n-2}"

"y'(x)=\sum_{n=0}^{\propto}a_n(r+n)(r+n-1)x^{r+n}+\sum_{n=0}^{\propto}a_n(r+n)x^{r+n}+\sum_{n=0}^{\propto}a_nx^{r+n+2}"

"a_0r^2x^r+a_1(r+1)^2x^{r+1}+\sum_{n=2}^{\propto}(a_n(r+n)^2a_{n-2})x^{r+n}=0"


The indicial equation is r2= 0, and hence r1 = r2 = 0. 

a1 = 0; the recurrence relation is

"a_n=\frac{a_{n-2}}{(r+n)^2}" ,n=2,3...

We conclude a1= a3 = a5 = … = 0, and since r = 0,

"a_2m=\frac{a_{2m-2}}{(2m)^2}" ,m=1,2...

"a_{2m}=\frac{(-1)^ma_0}{2^{2m}(m!)}" ,m=1,2...

"y_1(x)=a_0(1+\sum_{n=1}^{\propto}\frac{(-1)^ma_0}{2^{2m}(m!)})"

as y(1)=1, a0=1 so "y_1(x)=1+\sum_{n=1}^{\propto}\frac{(-1)^m}{2^{2m}(m!)}"

"J_0(x)=\sum_{n=1}^{\propto}\frac{(-1)^m}{2^{2m}(m!)}"


Since indicial equation has repeated roots, coeficients in second solution can be found using "a_n'(r)|_{r=0}"

"a_0(r)r^2x^r+a_1(r)(r+1)^2x^{r+1}+\sum_{n=2}^{\propto}(a_n(r)(r+n)^2a_{n-2}(r))x^{r+n}=0"

"a_1(r)=0, a_1'(0)=0"

"a_{n}(r)=\frac{a_{n-2}(r)}{r+n}" , n=2,3..

"a'_{2m+1}(0)=0" , m=1,2...

"a_{2m}(r)=\frac{a_{2m-2}(r)}{r+2m}= \frac{(-1)^ma_0}{(r+2)^2...(r+2m)^2}"

"\frac{a'_{2m}(r)}{a_{2m}(r)}=2(\frac{1}{r+2}+\frac{1}{r+4}...+\frac{1}{r+2m})"

"a'_{2m}(0)=2(\frac{1}{2}+\frac{1}{4}...+\frac{1}{2m})"

"H_m=\frac{1}{2}+\frac{1}{4}...+\frac{1}{2m}"

"y_2(x)=J_0(x)lnx1+\sum_{m=1}^{\propto}\frac{(-1)^{m+1}H_m}{2^{2m}(m!)}x^{2m},x>0"